### v1

# What Would an LLM Do? Evaluating Policymaking Capabilities of Large Language Models

Abstract
Large language models (LLMs) are increasingly being adopted in high-stakes domains. Their capacity to process vast amounts of unstructured data, explore flexible scenarios, and handle a diversity of contextual factors can make them uniquely suited to provide new insights for the complexity of social policymaking. This article evaluates whether LLMs’ are aligned with domain experts (and among themselves) to inform social policymaking on the subject of homelessness alleviation – a challenge affecting over 150 million people worldwide. We develop a novel benchmark comprised of decision scenarios with policy choices across four geographies (South Bend, USA; Barcelona, Spain; Johannesburg, South Africa; Macau SAR, China). The policies in scope are grounded in the conceptual framework of the Capability Approach for human development. We also present an automated pipeline that connects the benchmarked policies to an agent-based model, and we explore the social impact of the recommended policies through simulated social scenarios. The paper results reveal promising potential to leverage LLMs for social policy making. If responsible guardrails and contextual calibrations are introduced in collaboration with local domain experts, LLMs can provide humans with valuable insights, in the form of alternative policies at scale.
Introduction
Homelessness, defined by the United Nations (UN) as “the lack of a stable, safe, and adequate housing” (Special Rapporteur on the right to adequate housing 2022), is a growing crisis affecting cities and nations worldwide. According to the UN, an estimated 1.6 billion people globally lack adequate housing, with 150 million completely homeless. The OECD reports that in most developed countries, homelessness has sharply increased since 2022 (OECD 2024), with the U.S. alone recording more than 653,000 unhoused individuals in a single night in January 2023 (Daniel Soucy and Hall 2024). Nearly 15.6 % of the population of Johannesburg lived on the streets in 2022 (Statistics South Africa 2025), while Barcelona’s homeless population has grown by 61% between 2008 and 2022 (Xarxa d’Atenció a Persones Sense Llar (2023) XAPSLL). Homelessness represents a severe deprivation of basic human security and dignity, undermining sustainable development goals, and tackling its challenges is a core requirement for creating equitable, resilient societies where no one is left behind (United Nations 2015).
In response to these urgent social challenges, policymaking to alleviate homelessness should extend beyond material redistribution to confront a more profound need: ensuring unhoused individuals are treated as equals and fully integrated into the social fabric. This necessitates dismantling structural stigmatization and fostering genuine belonging—a task complicated by deeply entrenched societal biases, bureaucratic systems that often reduce people to “cases”, and the trauma-induced isolation experienced by those without stable shelter (Curto et al. 2025; Ranjit et al. 2024). As Wasserman and Clair (2009) underscore, homeless individuals frequently face “institutional humiliation” in shelters or services, exacerbating their marginalization. The World Bank’s Voices of the Poor further reveals that exclusion persists even after housing access, as societal rejection impedes employment, community participation, and psychological recovery (Narayan et al. 2000). Thus, effective homelessness mitigation policies must not only allocate resources but actively combat dehumanization, centering the agency and dignity of those affected, and recognize inclusion as both an ethical imperative and a pragmatic necessity.
Given the multifaceted nature of homelessness and the difficulty of anticipating policy outcomes in dynamic social systems, computational approaches offer a promising avenue for navigating this complexity and enable prospective policy testing. Recent advances in computational social science have demonstrated the potential of agent-based models (ABMs) to inform homelessness policymaking through dynamic modeling of complex socio-economic systems. Aguilera et al. (2025) developed an innovative framework for human development with ABMs, representing homelessness as a multidimensional deprivation, where agents interact within policy-sensitive environments. However, significant challenges remain in modeling human behavior within social simulations, such as the inability of capturing latent psychological factors like values, needs and emotional states (Dignum 2021), and the difficulty of developing generalizable models due to the dramatic differences in local policies, social networks, and resource availability across geographies (Aguilera et al. 2025). To address this limitation, there is active research on connecting ABMs with LLMs, demonstrating promising results in exploratory studies (Anthis 2025).
Simultaneously, we note that the increasing popularity of AI-driven tools like LLMs and their pervasive integration into societal systems make them increasingly likely to inform policy makers in various contexts, such as in finance (Febrian and Figueredo 2024) and law (Yun and Lee 2025). However, deploying LLMs in socially sensitive domains, such as homelessness policymaking, presents serious challenges. These include the amplification of societal biases (Agnew et al. 2024), generation of plausible but misleading outputs (“hallucinations”), a lack of genuine understanding of human suffering and social dynamics (Shanahan 2024), and inherent opacity that makes their reasoning difficult to audit. It is therefore imperative to proactively assess their risks and limitations in such high-stakes contexts.
To address these concerns, we introduce a benchmark with decision scenarios involving policy choices aimed at reducing homelessness, to assess the alignment of LLMs with domain experts in four different parts of the world (South Bend Indiana, USA; Barcelona, Spain; Johannesburg, South Africa; Macau SAR, China). The selection of policies included in the benchmark is grounded in the conceptual framework of the Capability Approach for human development (CA) (Sen 1999; Nussbaum 2011; Robeyns 2017). Unlike models that systematically prioritize the material needs of individuals (Maslow 1943), the CA considers opportunities that human beings have to lead a meaningful life with dignity (Sen 1999). From this perspective, homelessness can be considered as a deprivation of central capabilities (Nussbaum 2011). We perform an extensive empirical investigation around LLMs’ choices and judgments on the benchmark, including comparing them with those offered by domain experts in the geographical locations.
Finally, we propose a novel pipeline to automatically link policies with an ABM framework, intended to gauge the social impacts of LLM generated policies in a simulated social context. Specifically, this approach explores how policies inform agent behavior through an existing State-Action-Transition process (SAT) (Aguilera et al. 2024). Thus, the pipeline establishes a connection between algorithmic policy suggestion and simulated societal impact. Our work opens up new directions to leverage LLMs for supporting social policymaking in a manner that is scalable and non-invasive, where LLMs could be used to suggest new packages of policies aligned with the conceptual framework of the Capability Approach, and their impact could be examined via ABM simulations.
Related Work
The Capability Approach in Computational Frameworks
Although Sen and Nussbaum’s work on the Capability Approach (Sen 1999; Nussbaum 2011) constitutes a cornerstone in human development studies and the conceptual framework that underpins the Sustainable Development Goals, its operationalization in the practice of social policy making still has much potential to unveil (Robeyns 2017). From a computational social practice perspective, only recently has the CA been integrated into computational methods. There is an incipient corpus that aims to integrate the conceptual framework of the CA in agent-based modeling (Aguilera et al. 2025; Chávez-Juárez and Krishnakumar 2021), but to the best of our knowledge, our work constitutes the first application of CA as an audit framework for generative AI in social policy.
Evaluating LLMs for Social Policymaking
Recent research has explored the potential and limitations of LLMs in policymaking contexts. Jiao et al. (2025) systematically documented risks when deploying LLMs in high-stakes policy domains, highlighting issues of bias amplification and context blindness, which this paper aims to address. Work by Wei, Kumar, and Zhang (2025) provided evidence of how LLMs can perpetuate societal biases in resource allocation scenarios, raising critical questions about their suitability for welfare policymaking. The recurrent hallucinations in general problem solving (Xu, Jain, and Kankanhalli 2025) takes on particular significance in homelessness policymaking, where factual inaccuracies could directly impact the most vulnerable individuals.
While frameworks exist for evaluating LLM outputs in healthcare decision making (Kanithi et al. 2024) and in legal domains (Li et al. 2024), homelessness policy presents unique challenges due to its intersection with complex socioeconomic factors, stigma dynamics, and diverse cultural contexts. Our work presents a novel framework to specifically assess the ability of LLMs to inform policy making on the topic of homelessness alleviation.
Integrating LLMs with Agent-Based Social Simulations
The integration of LLMs with ABMs constitutes an emerging topic in computational social science (Chopra et al. 2024). Traditional ABMs have proved very helpful to inform social policy making in a non-invasive manner (Dignum 2021). LLMs have shown the potential to enrich the behavior of agents with reasoning-like abilities, memory, personality traits, and context sensitivity (Ferraro et al. 2025; Park et al. 2023; Horto 2023). Although conceptual frameworks have been presented to reliably connect LLMs with ABMs in socially sensitive frameworks (Ricci et al. 2024), the technical approaches still remain a challenge. In this work, we explore the potential impact of policies from our proposed benchmark in a simulated social context by adapting the behavior of agents accordingly in an agent-based modeling simulation context.
Methodology
In this section, we describe our methodology, an overview of which is illustrated in Figure 1.
Figure 1: Methodology overview: we construct a benchmark via a structured prompting strategy grounded in the Capability Approach, prompt various LLMs to act as policymakers, and analyze their choices through comparison with human expert judgments and a study of their projected societal impact using a modular agent-based modeling pipeline.
Benchmark Design
We present the components of a novel benchmark that aims to evaluate whether LLMs are aligned with domain experts, within the locations in scope, to inform homelessness alleviation policymaking [^1] The benchmark will be made publicly available later..
Dataset Composition
Locations and scope.
The benchmark covers four geographic contexts: Barcelona (Spain), Johannesburg (South Africa), South Bend (USA), Macau SAR (China), as well as a Universal context not tied to any particular city. The universal context is included to study the impact of context on choices. For each of the four specific locations, we generate 40 decision-making scenarios, while the universal category contains 10 scenarios. Each scenario frames a dilemma faced by a non–profit organisation (NPO) working to alleviate homelessness, e.g., how to allocate emergency housing funds, whether to prioritise medical outreach or job training, or how to intervene in the face of gentrification pressures. In total, the benchmark comprises 170 scenarios, each with 4 policy choices.
Scenario structure.
Every scenario is presented as a structured JSON object with four fields: Scenario (a title summarising the dilemma), Context (one or more paragraphs providing background), Policy_Options (an array of four policy proposals) and Main_capability_restoration (a list of capabilities targeted by each policy). The context is written in narrative form and spans at least 80 words, capturing locally relevant demographics, political debates, economic constraints and historical factors. Each policy option is described in a paragraph of at least 35 words and is annotated with one or more of Martha Nussbaum’s ten central human capabilities (see Appendix A) [^2] All Appendices are in the supplementary material.. These capabilities, summarized in the literature as universal freedoms that societies ought to protect, range from the ability to live a full life and access healthcare to the freedom to develop practical reason and to participate in political decision making (Nussbaum 2012). Using this normative framework ensures that each policy speaks to a specific aspect of human flourishing rather than merely enumerating material resources.
Scenario Generation
Baseline scenario design.
To seed the generation process, we first manually crafted a baseline scenario for Barcelona in consultation with domain experts. This scenario, rooted in current debates about housing scarcity (Xarxa d’Atenció a Persones Sense Llar (2023) XAPSLL), presented a homelessness dilemma alongside four plausible policy responses aiming to restore affected central capabilities. The manual scenario served both as a template for automated generation and as a qualitative anchor for calibrating LLM outputs.
Prompt-based generation.
Subsequent scenarios were generated using a frontier LLM (GPT–4.1) instructed to act as a public policy expert with knowledge on urban development and the Capability Approach. We provided the LLM with a detailed project description, a list of the ten central capabilities, guidance on scenario length and tone, and examples of desirable output. The model was then prompted to produce exactly 40 non–redundant scenarios for each city (and 10 for the universal category). Each prompt emphasized that scenarios must (1) present a localized public policy dilemma related to homelessness, (2) be either grounded in a documented real situation or be a plausible fictional event, (3) include context paragraphs of sufficient length, (4) propose four diverse and non–redundant policy options, and (5) annotate each policy with its primary capability restoration. To avoid hallucinated or extreme situations, we explicitly instructed the model not to invent unlikely events (e.g., large scale violence in low–crime areas) and to “think like a local expert who has studied homelessness for many years”. Generation was performed separately for each location to ensure contextual fidelity.
Quality control.
Generated scenarios were post–processed to remove redundancies and to correct factual inaccuracies. When the LLM produced incomplete or duplicate content, we regenerated the scenario with a revised prompt. We also verified that the distribution of capabilities across policy options was balanced so that no capability was systematically over-represented, and that the language was neutral and devoid of harmful stereotypes.
Policy Annotation and Ethical Grounding
Beyond the narrative context, every policy option in the benchmark is linked to one or more human capabilities. These annotations (explicitly linking to capabilities) serve two roles: (i) they can guide evaluators (human or model) toward understanding the ethical aspirations behind each intervention, and (ii) they provide a means for measuring whether LLM recommendations favour certain types of capabilities over others. Annotating the policies in the framework of the Capability Approch involved identifying the core goal of each policy—whether it primarily protects life and physical health, fosters social belonging, supports practical agency, or promotes other types of individual freedom—and mapping that to the relevant capability.
Human Expert Annotation
To create a gold-standard baseline against which LLMs can be compared, we engaged domain experts from three out of the four cities in the benchmark – Barcelona, Johannesburg, and South Bend. Each expert is a practitioner or scholar with contextualized knowhow on homelessness alleviation landscape in their city and works in close collaboration with local non–profits. We provided them with 10 out of the 40 scenarios relevant to their location as well as the 10 universal scenarios, asking them to prioritise the four policies in each scenario from 1 (most preferred) to 4 (least preferred). Experts were instructed that there was no single correct answer and that they should rank options based on their topic and local knowledge, as well as their ethical considerations. Through this process, we obtained three sets of human-derived rankings—one for each city—that serve as the benchmark’s reference recommendations.
Model Evaluation Tasks
While the benchmark could be used for various analyses, here we are interested in studying and evaluating how humans or machines make decisions as policymakers in two primary modes. In the top–choice task, they must select the single policy they would recommend; in the ranking task, they are asked to produce an ordered ranking of all four policy options. We undertake both tasks using several LLMs spanning different parameter scales, as described next.
LLM as a Policymaker
Our homelessness mitigation benchmark is presented to LLMs, which are prompted to respond as decision makers representing a non-profit organisation. Here we describe our choice of LLMs and baseline prompts.
Models.
We experiment with the following LLMs, listed in increasing order of the number of model parameters (which are shown in parentheses): Granite 3.1 8B Instruct (8 B), GPT 4.1 Mini (8 B), Mixtral 8X7B Instruct (around 47 B), Llama 3.3 70 B Instruct (70 B), Deepseek V3 (671 B), GPT 4.1 (estimated around 1.8 T). These models capture a broad spectrum of characteristics, spanning a range of sizes and architectures – from compact, efficiency-optimized models (e.g., Granite 3.1, GPT 4.1 Mini) to larger frontier models (e.g., Deepseek V3, GPT 4.1). This diversity enables comparisons across performance tiers, supporting robust conclusions about decision quality, consistency, and ethical alignment.
Prompts.
We prompt LLMs in two primary ways, one for each of the two tasks. In one approach, we ask the LLM to select one policy among the $4$ choices for scenarios in the benchmark. In another approach, we ask the LLM to provide an ordered ranking of their choices among the policies. In both cases, we also request the LLM to provide a brief justification for their response. For the GPT models, we employ a constrained reasoning prompt that guides models through four steps—summarizing the dilemma, identifying capabilities restored by each policy, assessing pros and cons, and outputting a ranking with justification. To encourage transparency without eliciting verbose chain–of–thought, we limit reasoning to under 150 characters per step. In addition, we conduct some experiments with ablation studies where LLMs are asked to make choices using additional considerations, such as paying heed to contextual information. We refer the reader to Appendix C where we provide some example prompts and prompt templates.
Assessing the Social Effects of Homelessness Mitigation Policies
We explore the social impact of the selected policies in an existing agent-model simulation (Aguilera et al. 2024) contextualized in the city of Barcelona, which is designed to evaluate the impact of homelessness alleviation policies.
Architecture.
In the pipeline, LLMs convert the natural language policy proposals they have previously selected into agents’ behavioral adjustments in the ABM. This is done by requesting the LLM to update the state-action-transition (SAT) matrix accordingly, which regulates the behavior of agents during the simulation. Due to space limitations, we relegate details about the mathematical foundations of the ABM pipeline to Appendix B. This approach enables us to test the effects of a policy in terms of needs satisfaction of the agents representing people experiencing homelessness (PEH).
Prompts.
We prompt an LLM by giving a specific role of a public policy expert and providing the following information: a description of the SAT matrix structure, the policy to be tested, a step-by-step description on how the matrix can be technically modified and the indication that the policies are framed in the Capability Approach. The LLM produces an update of the SAT matrix, modifying the behavior of agents in the simulation. More details on the specific prompt can be found in Appendix C.
Empirical Investigation
Refer to caption Refer to caption Refer to caption
(a) Top Choice Overlap (b) Justification Similarity (c) Ranking Comparison
Figure 2: Pairwise comparison of the choices and judgments of various LLMs using heat maps of the following: a) the fraction of common top choices, b) similarity between justifications for top choices (as measured by the ROUGE-L metric), and c) comparison of rankings of choices (as measured by the normalized Kendall tau distance).
Investigating LLMs’ Choices
Experiment.
We explore how the various LLMs compare in terms of their choices and judgments on the homelessness mitigation benchmark. Each LLM is asked to consider all $160$ contextualized decision scenarios – $40$ each across the $4$ locations – and make a judgment about their top choice and rankings. Figure 2 shows pairwise comparison plots between all $6$ LLMs under consideration in the form of heat maps, using different aggregate metrics across panels. Panel (a) displays the fraction of scenarios where the top choices for a pair of models are identical. Panel (b) shows the average similarity between the textual justifications of LLMs for their top choice, as measured by the ROUGE-L similarity metric. Panel (c) compares rankings of choices, as gauged by the normalized Kendall tau distance, which is $1$ for identical rankings and 1-1 for completely reversed rankings.
Results.
The plots in Figure 2 indicate that LLMs could potentially disagree about their top choices, particularly those from different model families. Panel (a) highlights that the smaller Granite 3.1 8B has less overlapping top choices, while other models choose the same policy around $70-80\%$ of the time. We also note that the ROUGE-L scores in panel (b) are substantially lower than the range of 0.40.50.4-0.5 (often regarded to represent moderate similarity), indicating that neutral prompts result in different justifications across LLMs on the benchmark. While there are well-known limitations of using ROUGE-L as a metric for comparing text, it remains popular for closely related tasks such as summarization. A comparison of rankings in panel (c) reflects similar trends. Overall, the results indicate that the choice of LLM affects policy choices and judgments on the benchmark.
Comparing Choices of LLMs vs. Experts
Experiment.
Our comparison with experts operates on two levels: agreement on the single top-ranked policy and similarity across the entire ranking of four policies, measured using the normalized Kendall tau distance. This allows us to assess not only the final choice but also the structure of preference. The comparison between the two human experts on Universal scenarios provides a baseline for inter-human agreement on decontextualized policy dilemmas.
Figure 3: Comparison of capabilities prioritized in the top policy choices of human experts and GPT-4.1 across all scenarios.
Table 1: Alignment b/w LLMs vs. experts: the low correlation for the South Bend pair, despite high top-choice agreement, suggests a divergence in underlying policy priorities.
Comparison Pair Top Choice Ranking
Agreement Corr. ($\tau$)
Location-Specific Scenarios
   GPT-4.1 vs. JHB Expert 50% 0.20
   GPT-4.1 vs. SB Expert 60% 0.10
Universal Scenarios
   GPT-4.1 vs. Exp. 1 (JHB) 20% 0.13
   GPT-4.1 vs. Exp. 2 (SB) 60% 0.33
   Human Baseline 40% 0.07
Results.
Our analysis reveals a moderate but nuanced alignment between LLMs and experts. As shown in Table 1, which compares experts with GPT-4.1, the degree of agreement varies by context. In Johannesburg, the LLM’s top choice matched the expert’s choice 50% of the time, with a weak ranking correlation (Kendall’s Tau) of 0.20. The South Bend case is particularly revealing: despite a higher top-choice agreement of 60%, the overall ranking correlation was a mere 0.10. This suggests that while the LLM and the South Bend expert could often agree on the single best policy, their underlying priorities for the remaining options diverged significantly.
The Universal scenarios provide the most telling baseline. The two human experts that annotate the universal scenarios agreed on a top policy in only 40% of cases, and their overall rankings showed almost no correlation ($\tau=0.07$). In this light, GPT-4.1’s alignment with the South Bend expert (60% agreement, $\tau=0.33$) is notable, as it exceeds the human-human baseline on both metrics. We conjecture that this is because the experts are taking contextual information into account while making their choice, which can be vital for effective policy making.
Figure 4: Comparing top choice overlap between 4 LLMs and an expert in Johannesburg, with and without prompting LLMs to pay attention to the context when choosing.
We investigate this further in Figure 4, where we consider scenarios in Johannesburg while comparing top choices of the four non-GPT LLMs with those from the expert in four ways: over the 10 contextualized (Local) scenarios and the 10 Universal scenarios, both with and without additional prompting to the LLM to explicitly consider the local geographical context. We observe that in general: 1) there is more expert choice overlap in Local rather than Universal settings, and 2) specifying the context impacts LLMs’ choices and yields more overlap, even when scenarios are Local, i.e., already somewhat contextualized by design.
Beyond pure agreement rates, we find systematic divergence in the types of capabilities prioritized, as visualized in Figure 3. Experts’ top choices frequently centered on policies restoring practical reason (e.g., skills training, participatory governance) and affiliation (e.g., peer support, community integration). In contrast, GPT-4.1 demonstrated a stronger tendency to select policies that protect against immediate physical threats, prioritizing life, bodily health, and bodily integrity. The LLM appears to exhibit a form of heightened risk aversion focused on fundamental safety, whereas experts more readily balance safety with long-term empowerment and social cohesion. For instance, in the scenario titled “Safe-Parking for Vehicle Dwellers” (South Bend), the human expert preferred ‘Conversion-to-Housing Incentive Grants‘ to transition residents into stable rooms, emphasizing long-term capability expansion. GPT-4.1 favored a ‘Monitored Safe-Parking Lot‘, echoing its bias toward scalable, immediate hazard reduction.
This value divergence is starkly illustrated in a Johannesburg scenario regarding an open-air drug scene. The expert prioritized a “Broken Windows” policing surge (Capability: bodily integrity via public order). GPT-4.1 ranked this last, instead choosing a supervised consumption site (Capabilities: life, bodily health). Its verbal justification noted that policing “risks rights for limited gain,” while the site “saves most lives”. This highlights a fundamental difference: the expert chose an enforcement-based solution to restore order, while the LLM opted for a health-based, harm-reduction solution to prevent death. In summary, while an LLM’s choices could potentially be as consistent with an expert as that expert is with another, its underlying decision-making calculus suggests a discernible bias towards scalable, safety-first interventions over more relational or structural ones.
Gauging the Impact of Policies using ABMs
Table 2: Comparison of of how policies recommended by LLMs and experts fulfill the needs of PEH agents in the simulation.
Scenario Physiological Safety Belonging Self-esteem Mean Std. p-value Mean Std. p-value Mean Std. p-value Mean Std. p-value Scenario 1    LLM Policy +0.019 -0.005 0.001 +0.032 -0.023 0.001 +0.021 -0.008 0.030 +0.037 -0.025 0.001    Expert Policy -0.004 -0.004 0.457 +0.026 -0.023 0.006 -0.036 -0.011 0.009 +0.027 -0.025 0.015 Scenario 3    LLM Policy +0.012 -0.010 0.080 +0.031 -0.025 0.001 +0.012 -0.013 0.302 +0.036 -0.027 0.001    Expert Policy +0.011 -0.011 0.029 +0.030 -0.025 0.001 +0.009 -0.017 0.092 +0.035 -0.028 0.001 Scenario 5    LLM Policy +0.014 -0.020 0.134 +0.031 -0.026 0.002 +0.014 -0.034 0.322 +0.038 -0.030 0.002    Expert Policy +0.002 -0.020 0.909 +0.029 -0.025 0.002 -0.016 -0.032 0.036 +0.033 -0.029 0.003
Experiment.
We evaluate the concrete impacts of some of the benchmarked policies in terms of whether they satisfy the needs of the agents (representing PEH) as described in Section Assessing the Social Effects of Homelessness Mitigation Policies. We run simulations with 80 agents for 1450 steps (corresponding to a time frame of 2 months in the model), and use Deepseekr1 (with a temperature of 0.1) to update the SAT matrix on the scenarios where the expert top policy prioritization choice differs from the GPT4.1 top policy prioritization for the city of Barcelona (these are scenarios 1, 3 and 5 in the benchmark). The ABM framework used in this experiment is specifically designed for the city of Barcelona at the moment and will be extended further in future work (please see Appendix B for details). We run 10 simulations without the behavioral changes from the applied policy (with the base SAT matrix of the model) and 10 simulations with the behavioral changes induced by the policy (with the LLM-updated SAT matrix). This approach provides us with insights on how the policies selected by the expert and the LLMs differ in terms of the satisfaction of the needs of the agents (representing PEH).
Results.
Table 2 summarizes the changes observed in the ABM simulation when applying either the LLM or the expert policy. For each category of needs, ‘Mean’ represents the difference of the average satisfaction value for PEH agents with or without the policy. ‘Std’ represents the difference in the dispersion of needs satisfaction with or without policies. And finally, ‘p-value’ is associated with a standard t-test to assess whether or not the changes in terms of needs’ satisfaction were statistically significant.
Based on the results of Table 2, the LLM-recommended policies seem to demonstrate superior overall balance compared to the expert-recommended policies. While both approaches notably enhanced safety and self-esteem needs, the improvements from LLM-implemented policies were more substantial. Additionally, LLM implementations consistently prevented the negative effects on belonging needs that were seen with expert policies in scenarios 1 and 5. Both policy types successfully reduced outcome variability (negative change in std values), but LLMs achieved this while maintaining positive or neutral effects across all need categories—a balance experts couldn’t sustain. More detailed results are available in Appendix D.
Discussion and Conclusions
This paper introduces a novel benchmark to evaluate whether LLMs are aligned with human domain experts to inform social policymaking, in particular on the topic of homelessness alleviation. The benchmark includes 170 homelessness-related policy scenarios, grounded in the Capability Approach for human development, across four geographic contexts. It further presents an automated pipeline that links LLM-generated policy proposals to an agent-based modelling framework. This enables to explore the social impact of the recommended policies in a non-invasive manner through simulated social scenarios. The article opens a path to leverage LLMs for social policy making, providing humans with alternative, contextualized policies at scale, and an indication of their social impact.
The empirical investigation unveils that LLMs policy recommendation patterns seem to apply a highly stable internal heuristic, i.e. prioritize immediate physical safety and broad coverage—across different contexts. While this leads to dependable internal logic, it can also expose a context-blind rigidity. Experts, in contrast, demonstrably tailored their decisions to address locally encoded sociopolitical realities that can aggravate the ostracism faced by PEHs, such as related to ethnicity (in South Africa) or to religious minorities (in South Bend). Moreover, our analysis reveals a moderate higher alignment on LLMs’ top choice with the two experts located in the Global North. However, when asking LLMs to focus in a contextualized scenario, the results become more closely aligned with the local expert (as we describe for Johannesburg). Our findings also compare the social impact of the policies selected by LLMs vs those selected by the domain expert in the city of Barcelona, in an agent-based modelling framework. The results obtained in the simulated social scenario show that LLMs recommended policies could be better fulfilling the overall needs of PEH population.
As in most computational approaches to complex social topics, the paper contains some limitations that will be addressed in future work. First of all, policies were manually annotated by only one domain expert per location. Although these domain experts work in close collaboration with specialized local non-profits on homelessness alleviation, feedback from additional experts will provide a broader representation of the local context. Secondly, our study has only been conducted with LLMs in English, which is not the most spoken language in all the locations in scope. Multilingual analysis will be included in future analysis. Finally, the agent-based modelling framework only handles scenarios for one of the locations in scope (Barcelona) and is only able to simulate some of the policies in scope. Future work will expand the simulated scenarios to obtain more comprehensive information on the social impact of the proposed policies.
These results reveal promising potential for LLMs to provide new insights on social policy making, if the analysis is rooted in a conceptual ethical framework which is relevant to the social topic in scope – our study is framed in the Capability Approach for human development – and when responsible guardrails and contextual calibrations are introduced in collaboration with local domain experts.
Appendix A Nussbaum’s Central Human Capabilities
We list and describe the 10 central human capabilities in the Capability Approach, based on (Nussbaum 2011).
  • Life : Being able to live to the end of a human life of normal length; not dying prematurely, or before one’s life is so reduced as to be not worth living.
  • Bodily Health : Being able to have good health, including reproductive health; to be adequately nourished; to have adequate shelter.
  • Bodily Integrity : Being able to move freely from place to place; to be secure against violent assault, including sexual assault and domestic violence; having opportunities for sexual satisfaction and for choice in matters of reproduction.
  • Senses, Imagination and Thought : Being able to use the senses, to imagine, think, and reason—and to do these things in a ”truly human” way, a way informed and cultivated by an adequate education, including, but by no means limited to, literacy and basic mathematical and scientific training. Being able to use imagination and thought in connection with experiencing and producing works and events of one’s own choice, religious, literary, musical, and so forth. Being able to use one’s mind in ways protected by guarantees of freedom of expression with respect to both political and artistic speech, and freedom of religious exercise. Being able to have pleasurable experiences and to avoid non-beneficial pain.
  • Emotions : Being able to have attachments to things and people outside ourselves; to love those who love and care for us, to grieve at their absence; in general, to love, to grieve, to experience longing, gratitude, and justified anger. Not having one’s emotional development blighted by fear and anxiety. (Supporting this capability means supporting forms of human association that can be shown to be crucial in their development.)
  • Practical Reason : Being able to form a conception of the good and to engage in critical reflection about the planning of one’s life. (This entails protection for the liberty of conscience and religious observance.)
  • Affiliation : Being able to live with and toward others, to recognize and show concern for other humans, to engage in various forms of social interaction; to be able to imagine the situation of another. (Protecting this capability means protecting institutions that constitute and nourish such forms of affiliation, and also protecting the freedom of assembly and political speech.) Having the social bases of self-respect and non-humiliation; being able to be treated as a dignified being whose worth is equal to that of others. This entails provisions of non-discrimination on the basis of race, sex, sexual orientation, ethnicity, caste, religion, national origin and species.
  • Other Species : Being able to live with concern for and in relation to animals, plants, and the world of nature.
  • Play : Being able to laugh, to play, to enjoy recreational activities.
  • Control Over One’s Environment : Being able to participate effectively in political choices that govern one’s life; having the right of political participation, protections of free speech and association. Material. Being able to hold property (both land and movable goods), and having property rights on an equal basis with others; having the right to seek employment on an equal basis with others; having the freedom from unwarranted search and seizure. In work, being able to work as a human, exercising practical reason and entering into meaningful relationships of mutual recognition with other workers.
Appendix B Urban ABM Backbone: Theoretical Foundations
This ABM framework comes from the work of (Aguilera et al. 2024). It relies on the set of Maslow’s need categories ( Physiological, Safety, Belonging, Esteem) as $C$ and the set of needs within each category $c\in C$ as $N_{c}$ (Physiological: food, shelter, sleep, health; Safety: clothing, financial security, employment, education; Belonging: family, friendship, intimacy; Esteem: freedom, status, self-esteem) thus resulting in 14 needs across each category. Each agent $a_{i}\in\mathcal{A}$ maintains:
  • A status $s\in\{\texttt{homeless},\texttt{employed},\texttt{unemployed}\}$
  • A need-satisfaction vector $\mathbf{n}^{(t)}\in[0,1]^{14}$ at timestep $t$
  • A satisfaction-Action Transition (SAT) matrix based on each status $\mathbf{M}_{a}\in\mathbb{R}^{14\times 11}$. This SAT matrix maps the 11 possible actions (See Appendix) that agents can take depending on status to a need satisfaction bonus.
An importance function maps every need category to the weight that the agent assigns to it, $\operatorname{Imp}:C\to[0,1]$. At time-step $t$, the need satisfaction level of need $n\in N_{c}$ (for some $c\in C$) is given by the need satisfaction level function $\operatorname{NSL}_{t}:\bigcup_{c\in C}N_{c}\to[0,1]$, which maps every need (across all categories) to its current degree of fulfilment. It can be written as an iterative function:
$\operatorname{NSL}_{t}(n)$ $=\gamma_{n,s(n)}\cdot\operatorname{NSL}_{t-1}(n)$
where $\gamma_{n,s}\in[0,1]$ is the decay rate for a need $n$ and an agent with status $s$. The customization of the decay rates based on the nature of need and the agent’s status $s$ allows to create a differentiation between the personal circumstances of the agents.
Agent decision-making then follows a greedy algorithm. At each timestep t=1\ t=1 hour, agent $a$ with status $s$:
  1. 1.
    Identifies dominant unmet need: $n^{*}=\operatorname*{arg\,min}_{j}\,\mathbf{n}_{j}^{(t)}$
  2. 2.
    Selects action $\alpha$ via:
    $\alpha_{t}=\underset{\alpha\in A}{\mathrm{arg}\,\mathrm{max}}\left[\sum_{c\in C}\left(\sum_{\mathclap{n\in N_{c}}}M_{a}[n]\cdot(1-NSL_{t}(n)\right)\cdot Imp(c)\right].$ (3)
  3. 3.
    Updates needs:
    $\mathbf{n}_{j}^{(t+1)}=\gamma_{{n}_{j},s}\mathbf{n}_{j}^{(t)}+\delta_{j,n^{*}}\mathbf{M}_{a}[n^{*},\alpha]$
    with $\delta_{j,n^{*}}$ the Kronecker delta symbol :
    $\delta_{j,n^{*}}=\begin{cases}1&\text{if }j=n^{*}\\ 0&\text{if }j\neq n^{*}\end{cases}$
Thus the SAT Matrix is the key that is used by the agents to inform their behaviors by selecting the action to accomplish at each step of the simulation.
LLM-ABM Integration
The critical innovation that we bring to address the problem of scalability for policy testing lies in bridging social policy narratives with expected behavioral adjustments through LLM translation.
Indeed usual policy testing using ABM through SAT matrices faces two fundamental limitations:
$\underbrace{\mathcal{P}_{\textrm{narrative}}}_{\begin{subarray}{c}\text{Natural language}\\ \text{policy proposals}\end{subarray}}\xrightarrow{\textit{manual}}\underbrace{\Delta\mathbf{M}}_{\begin{subarray}{c}\text{SAT matrix}\\ \text{adjustments}\end{subarray}}\quad\text{is}\quad\begin{cases}\text{work intensive}\\ \text{opaque in value alignment}\end{cases}$ (1)
While informing the behavior of agents by calling LLMs for each decision of each agent during a simulation can be a technical solution instead of relying on a SAT matrix, this approach appears to be constrained by scalability concerns and limited number of tokens context for LLMs. Indeed, numerous LLMs call to simulate the interaction between agents result in a large amount of tokens to be processed and generated by LLMs, which in returns make the simulation costly in terms of money (if you need an access to a private API), energy consumption, and time (Feng et al. 2025).
Our pipeline addresses this by formalizing LLMs as differential policy operators:
$\mathcal{P}_{\textrm{narrative}}\xrightarrow{\textit{LLM}_{\theta}}\Delta\mathbf{M}\quad\textrm{s.t.}\quad$ $\|\Delta\mathbf{M}\|_{\infty}\leq 0.03$
where $\theta$ denotes ethical priors (capability approach) encoded in the LLM’s prompt (See appendix).
Causal Interpretability Framework
The integration establishes a closed-loop causal pathway for ethical impact assessment:
$\underbrace{\mathcal{P}}_{\text{Policy}}\xrightarrow{\textit{LLM}}\underbrace{\Delta\mathbf{M}}_{\text{Adjustment}}\xrightarrow{\textit{ABM}}\underbrace{\Delta\mathbb{E}[\alpha]}_{\text{Behavior}}\xrightarrow{}\underbrace{\Delta\mathbb{E}\left[\mathbf{n}^{(T)}\mid s=\texttt{homeless}\right]}_{\text{Equity Impact}}$ (2)
Policies $\mathcal{P}$ thus inject perturbations via precondition (e.g., $s=\texttt{homeless}$) and SAT overrides:
$\mathbf{M}_{a}\leftarrow\begin{cases}\mathbf{M}_{a}+\Delta\mathbf{M}&\text{if }\phi(a)=\texttt{True}\\ \mathbf{M}_{a}&\text{otherwise}\end{cases}$ (3)
enabling targeted assessment of distributional effects while preserving mechanistic accountability through the ABM’s core. It shows that the policies can used to target other groups of people than just people experiencing homelessness, thus expanding the potential usage of this framework.
Appendix C Prompts for LLMs
Prompts for LLM Responses for the Benchmark
Baseline prompt for top choices.
Below, we provide an example prompt for requesting the LLM to choose from among the policies in a decision scenario in the homelessness mitigation benchmark. This example is for the first decision scenario in Barcelona from the benchmark.
1Task:
2Act as an expert policymaker. Your goal is to study the decision scenario and to propose your top policy among the policy choices.
3
4Instructions:
51. Decision Scenario Analysis:
6 - Analyze the context of the decision scenario as well as the potential policy choices for helping alleviate homelessness.
7
82. Response Structure:
9 - Answer: Start with the policy choice number wrapped in squared brackets (e.g., [1], [2], [3], [4]) based on your analysis.
10 - Justification: Briefly explain why you chose the selected option.
11
12**Input Format:**
13
14Scenario:
15{ Insert the scenario name and context here }
16
17Policy Options:
18
19 {Insert the list of policy options along with their descriptions }
20
21---
22Your Turn:
23
24Scenario:
25Title: Deploying EU Recovery Funds: Temporary Modular Housing vs. Long-Term Solutions
26Context: Barcelona has received 24 million in NextGenerationEU recovery funds earmarked for urgent housing innovation. City planners propose using part of the money to expand APROPBarcelonas experimental shipping-container micro-apartment programso that unsheltered residents near Plaça de les Glòries can move into 92 prefabricated units within eighteen months. Critics argue the funds should instead reinforce permanent Housing First placements or bolster rent-supplement vouchers in adjacent municipalities where rents are lower. Local neighborhood groups worry that another modular block will accelerate gentrification pressures and erase remaining community gardens created on the former elevated ring-road site. With municipal elections a year away, the non-profit Habitat Digne BCN must pick one strategic path to recommend before City Hall freezes the allocation.
27
28Policy Options:
291: Scale Up APROP to 92 Additional Units at Glòries
30Dedicate the full 24 million to fast-track construction of stackable, energy-efficient container micro-homes on municipally owned land beside the new tramway link. Each 30 m2 unit includes private bathrooms, kitchenette, and climate control. A social-work team would provide on-site case management for up to five years, prioritizing rough sleepers with chronic illnesses, thereby stabilizing residents health and safety while the districts formal social-housing queue progresses.
312: Bolster Permanent Housing First Placements across AMB
32Redirect two-thirds of the grant toward purchasing existing flats in low-vacancy municipalities like Santa Coloma and Sant Adrià, integrating tenants directly into mainstream neighborhoods with individualized support. The remaining funds seed a revolving rent-guarantee pool that reassures private landlords, empowering participants to sign standard leases and exercise normal tenancy rights rather than residing in temporary structures.
333: Rent-Supplement Vouchers Plus Job-Link Services
34Create 600 time-limited housing vouchers pegged to 40 % of median rent, valid city-wide but portable across metropolitan transit zones. Pair vouchers with tailored employment coaching at Barcelona Activa hubs, aiming to lift recipients above the poverty threshold within two years. Although less capital-intensive, this path assumes the private market can absorb increased demand without driving rents.
354: Community Land Trust & Garden Preservation Compact
36Invest the funds in purchasing the contested Glòries plots for a resident-led community land trust that mixes 50 permanent cooperative apartments with ground-floor urban farms and cultural studios. This hybrid model keeps green commons intact, promotes ecological stewardship of pollinator gardens, and embeds arts-based therapy workshops, supporting holistic well-being for formerly unhoused residents while resisting speculative real-estate cycles.
37
38
39Respond using the structured format as described above.
Variation for ranking.
In a variation, we request the LLM to provide a ranking among the choices in a decision scenario. The only change in the prompt lies in our primary instruction to the LLM for the task, which is as follows: “Act as an expert policymaker. Your goal is to study the decision scenario and to propose an ordered ranking of the provided policy choices, from most preferred to least preferred. ”
Variation for contextual emphasis.
In a variation, we request the LLMs to pay heed to the context as they make their choice. The only change in the prompt lies in our primary instruction to the LLM for the task, where we append some text. An example of the instruction (for Johannesburg) is as follows: “Act as an expert policymaker. Your goal is to study the decision scenario and to propose your top policy among the policy choices. This decision scenario is set in the city of Johannesburg in South Africa. Please try to take the location into account for your analysis.”
ABM Pipeline Prompt
Here we share our prompt for using LLMs to connect policies to social impact using ABMs.
1You are a public policy expert with deep knowledge of social justice, urban development,
2and the CAPABILITY APPROACH developed by Martha Nussbaum.
3You specialize in designing realistic and ethically sensitive agents-based MODELS for use in simulated environments involving large language models (LLMs).
4
5Here you are given a SAT matrix mapping needs to actions for people experiencing homelessness.
6
7MATRIX STRUCTURE EXPLANATION:
8-The matrix has 14 rows (needs) and 11 columns (actions)
9- matrix[row][column] = matrix[need_index][action_index] = satisfaction_value
10- Row 0 = food, Row 1 = shelter, Row 2 = sleep, etc.
11- Column 0 = go_grocery, Column 9 = go_reception_center, etc.
12
13{
14 actions: [go_grocery, go_hospital, go_shopping, go_leisure, invest_education, sleep_street, beg, steal_food, steal_clothes, go_reception_center, go_prison],
15 needs: [food, shelter, sleep, health, clothing, financial security, employment, education, family, friendship, intimacy, freedom, status, self-esteem],
16 matrix: [
17 [1.0, 0.0, 0.0, 0.4, 0.0, 0.0, 0.15, 0.7, 0.0, 0.5, 0.0],
18 [0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.7, 0.0],
19 [0.0, 0.0, 0.0, 0.0, 0.0, 0.7, 0.0, 0.0, 0.0, 0.0, 0.0],
20 [0.3, 1.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0],
21 [0.0, 0.0, 1.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.8, 0.0, 0.0],
22 [0.0, 0.0, 0.0, 0.0, 0.5, 0.0, 0.5, 0.0, 0.0, 0.0, 0.0],
23 [0.0, 0.0, 0.0, 0.0, 0.5, 0.0, 0.5, 0.0, 0.0, 0.0, 0.0],
24 [0.0, 0.0, 0.0, 0.0, 1.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0],
25 [0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0],
26 [0.0, 0.0, 0.0, 0.4, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0, 0.0],
27 [0.0, 0.0, 0.0, 0.1, 0.0, 0.0, 0.0, 0.0, 0.0, 0.3, 0.0],
28 [0.0, 0.0, 0.0, 0.3, 0.4, 0.0, 0.4, 0.0, 0.0, 0.0, 0.0],
29 [0.0, 0.0, 0.0, 0.7, 0.6, 0.0, 0.6, 0.0, 0.0, 0.0, 0.0],
30 [0.0, 0.0, 0.6, 0.5, 0.0, 0.0, 0.0, 0.0, 0.6, 0.0, 0.0]
31 ]
32}
33
34HERE IS THE POLICY YOU MUST CONSIDER:
35{policy}
36
37YOUR TASK IS TO READ THE POLICY CAREFULLY AND MAKE CHANGES IN THE SAT MATRIX BASED ON THE POLICY.
38
39MATRIX MODIFICATION RULES:
40- You MUST modify the exact matrix positions that correspond to the policy
41- Use the correct row index (need) and column index (action)
42- Changes must be realistic: typically 0.01 to 0.03 for existing values, max 0.02 for zero values
43- You MUST make at least 2-4 specific changes
44- Focus on the most relevant need-action combinations
45- CRITICAL: NEVER add or remove rows/columns - keep exactly 14 rows and 11 columns
46- CRITICAL: Only modify VALUES within existing matrix positions
47
48CRITICAL: When you modify matrix[row][column], ensure:
49- Row index matches the need you want to affect (0 =food, 1 =shelter, 2 =sleep, 3 =health, etc.)
50- Column index matches the action you want to affect (0 =go_grocery,9 =go_reception_center, etc.)
51
52STEP-BY-STEP PROCESS:
531. Identify which ACTIONS are most affected by the policy
542. Identify which NEEDS are most improved by the policy
553. Find the exact matrix[need_index][action_index] positions
564. Make small realistic changes (0.01-0.03)
575. Verify your changes make logical sense
58
59IMPORTANT: YOU MUST MAKE AT LEAST 2-4 SPECIFIC CHANGES to matrix values. Do not return the original matrix unchanged.
60
61CRITICAL JSON FORMAT REQUIREMENTS:
62- Your response MUST contain valid JSON only - NO COMMENTS in the JSON
[…]
64- Do NOT add explanatory text inside the JSON structure
65- The JSON must be complete and properly closed
66- Use exactly this format:
67
68json
69{
70 actions: [go_grocery, go_hospital, go_shopping, go_leisure, invest_education, sleep_street, beg, steal_food, steal_clothes, go_reception_center, go_prison],
71 needs: [food, shelter, sleep, health, clothing, financial security, employment, education, family, friendship, intimacy, freedom, status, self-esteem],
72 matrix: [ ]
73}
74
75AFTER the JSON, you may provide your reasoning and explanation in plain text.
76
77CAPABILITY LIST:
781. Life
792. Bodily Health
803. Bodily Integrity
814. Senses, Imagination, and Thought
825. Emotions
836. Practical Reason
847. Affiliation
858. Other Species
869. Play
8710. Control over Ones Environment
88
89CONTEXT:
90
911. Introduction
92
93As large language models (LLMs) become more integrated into public services and civic decision-support tools, understanding how these models perform in ethically sensitive, real-world contexts is critical. Local governments and non-profit organizations often face complex social policy decisions, such as allocating limited housing, food, or job-training resources. These choices often involve nuanced ethical considerations, contextual responsiveness, and the need for consistent, rational judgment.
94
95Although LLMs are increasingly capable in generating plausible, articulate responses, their actual alignment with human ethical norms and local contextual factors in decision-making has not been sufficiently explored or benchmarked in the domain of social good. This is particularly relevant as LLMs continue to be deployed more broadly, including potentially playing the role of decision-makers in agent-based models for situations with public policy implications. This study seeks to address this gap.
96
97In particular, we adopt Nussbaums Capability Approach as a guiding framework for both scenario design and evaluation, recognizing that human development entails more than resource distributionit concerns restoring and expanding peoples substantive freedoms to live lives they have reason to value. This philosophical lens allows us to examine not only what LLMs decide, but what capabilities they prioritize in simulated public dilemmas.
98
992. Objectives
100
101This project aims to evaluate how large language models simulate public decision-making in social good contexts, focusing on:
102
103 Ethical alignment: Do model-generated decisions reflect fairness, equity, and harm reduction?
104 Capability sensitivity: Do models identify and reason around the restoration or expansion of key human capabilities, as defined in the Capability Approach?
105 Contextual awareness: Do models respond appropriately to local factors and stakeholder needs?
106 Consistency: Are model responses stable across similar or evolving prompts?
107 Value trade-off sensitivity: Does the model recognize competing values (e.g., efficiency vs. equity) and reflect moral pluralism?
108
109[Note that these are illustrative dimensions for evaluation and should be re-assessed. We should choose the evaluation criteria based on requirements around the ABMs.]
110
1113 Scenario Design
112
113Each scenario is framed from the perspective of an NPO leader or board, who must select among multiple feasible interventions given limited resources and organizational mission. The intent is to capture the ethical, practical, and contextual complexity facing NPOs in everyday operations: balancing immediate needs against long-term development, honoring the dignity of service users, and making transparent value trade-offs.
114
115To anchor these choices in a robust normative framework, every policy option will be explicitly annotated with its primary restoration of one or more of Martha Nussbaums Central Human Capabilities. The full list is as follows:
1161. Life.
1172. Bodily Health.
1183. Bodily Integrity.
1194. Senses, Imagination, and Thought.
1205. Emotions.
1216. Practical Reason.
1227. Affiliation.
1238. Other Species.
1249. Play.
12510. Control over Ones Environment.
126
127
1283.1 Implementation Framework
129
130 Agents are individual entities with a status label (e.g., homeless, employed, student). Each agent keeps a 14-element Need-Satisfaction Level (NSL) vector ranging from 0.0 - 1.0 that corresponds to the needs order in the SAT matrix.
131 Each simulation tick represents one in-game hour. An agent:
132 - Identifies its currently most pressing need (minimum NSL).
133 - Looks up every action available to its status in the SAT matrix and retrieves the satisfaction coefficient for that need.
134 - Calculates expected utility = 0.7 × SAT coefficient (the 0.7 mimics diminishing real-world returns).
135 - Selects the action with the highest expected utility.
136 After performing the action the agent updates its NSL for the satisfied need: NSL[n] = min(NSL[n] + 0.7 × SAT[n,action], 1.0)‘. All other needs decay slightly each hour to simulate ongoing deprivation.
137 Policies are injected as Norm objects that:
1381. Optionally override the chosen action ({agx.chosenaction = go_reception_center’}, etc.).
1392. Add a small delta (<= 0.03), usually <5% to specific SAT matrix cells for homeless agents, thereby locally boosting how much a given action fulfills a target need.
140 Because policies only tweak a handful of cells and respect hard caps, aggregate NSL values change gradually and never jump to 1.0 instantlythis yields realistic policy impact curves over multi-day simulations.
141 During batch experiments we run N iterations with different random seeds for a baseline (no policy) and for the policy scenario. At the final step of each run we aggregate NSL values across all homeless agents, producing per-run means that are then compared with t-tests.
142 The framework therefore provides a transparent, auditable mapping: <Policy text> <SAT tweaks> <Agent behaviour> <Population-level outcomes>.
143
144
1454. Expected Outcomes
146
147 A publicly available benchmark dataset of policy-style scenarios for evaluating LLM reasoning in social-good contexts.
148 A validated rubric for assessing ethical alignment, pluralistic reasoning, and policy awareness in LLMs.
149 A peer-reviewed paper reporting empirical findings and proposing best practices for LLM use in civic settings.
150 Guidance for public agencies and nonprofits exploring AI assistance in planning and outreach efforts.
151 An automated pipeline taking the policy and implementing it in this framework (THIS IS THE WORK YOU ARE HELPING TO DO).
152
153
154RESPONSE FORMAT:
1551. First, provide the complete updated SAT matrix in valid JSON format (no comments inside JSON)
1562. Then, explain your reasoning for the changes you made
157
158RETURN THE UPDATED SAT MATRIX IN THE EXACT JSON FORMAT SHOWN ABOVE - NO COMMENTS INSIDE THE JSON!
Appendix D Additional Experiments
In this section, we include some experimental results that we could not include in the main file due to space constraints.
Comparing LLMs’ Choices
Refer to caption Refer to caption Refer to caption
(a) Barcelona (b) Johannesburg (c) South Bend
Figure 5: Pairwise comparison of the top choices of various LLMs (with each other as well as a local expert) using heat maps for the three geographic regions with expert assessments. 10 contextualized scenarios are used while comparing experts with LLMs, whereas all 40 contextualized scenarios are considered while comparing LLMs with each other.
We conduct additional experiments regarding location-specific top choices by LLMs on the benchmark. Figure 5 compares top choices by experts and LLMs for each of the three locations assessed by experts – Barcelona, Johannesburg, and South Bend. Only the 10 location-specific contextualized scenarios that were assessed by experts are considered when LLMs are compared with experts, but all 40 contextualized scenarios are considered for comparisons between LLMs. The plots provide additional detail around some of the results outlined in the main file. Note that LLM top choices are more similar to each other and the expert in the locations of Barcelona and South Bend than in Johannesburg.
ABM Policy Comparison
Table 3: Policy Comparison across Scenarios
Policy Physiological Safety Belonging Self-esteem
Mean Std. Dev. Mean Std. Dev. Mean Std. Dev. Mean Std. Dev.
Scenario 1
No policy 0.786 0.013 0.957 0.025 0.821 0.024 0.939 0.028
LLM policy 0.805 0.008 0.988 0.002 0.842 0.015 0.977 0.003
Expert policy 0.782 0.009 0.983 0.002 0.785 0.012 0.966 0.004
Scenario 3
No policy 0.790 0.019 0.957 0.026 0.830 0.030 0.940 0.030
LLM policy 0.802 0.009 0.988 0.002 0.842 0.017 0.976 0.003
Expert policy 0.801 0.008 0.987 0.001 0.839 0.013 0.975 0.002
Scenario 5
No policy 0.785 0.027 0.956 0.027 0.819 0.044 0.937 0.032
LLM policy 0.798 0.007 0.987 0.001 0.833 0.010 0.975 0.002
Expert policy 0.787 0.007 0.985 0.002 0.803 0.012 0.970 0.003
In this table, we detail the results that are presented in Section 4.3 in terms of needs category for the agents representing PEH per scenario. These results were obtained by modifyng the SAT matrix by calling deepseekr1 API with a temperature of 0.1 and using the prompt presented in C.